OpenStudy (anonymous):

A soccer player kicks with the mass of 0.42 KG. The ball travels at a velocity of 28m/s for 1.6 seconds until the goalie stops it. What is the average acceleration of the soccer ball?

OpenStudy (anonymous):

I got 16.8

OpenStudy (kainui):

16.8 Volts per Angstrom?

OpenStudy (anonymous):

No.. 16.8 for the acceleration

OpenStudy (anonymous):

Is my answer wrong? I did 28*1.6-28

OpenStudy (kainui):

I think it is, why did you do the calculation like that?

OpenStudy (kainui):

I think it is wrong*

OpenStudy (anonymous):

Hmm, how would you have solved the question than?

OpenStudy (anonymous):

Because the formula for acceleration is Final Velocity-Initial Velocity..

OpenStudy (kainui):

I suggest keeping your units with you when you do a calculation, it will give you ungodly answer checking abilities. That is almost right, but let's think about it. Acceleration is a change in velocity, so you've got that going for you which is good. However, from my interpretation of the question, the ball's velocity is relatively constant from the point it was kicked to the point it was stopped. So we're really thinking about an impulse force stopping it from the goalie where the velocity abruptly stops when his foot hits it and very little deceleration happens over the course of a little over a second. So as you've calculated it, this is what your units look like: \[28\frac{ m }{ s }*1.6 s-28\frac{ m }{ s } \neq ???\frac{ m }{ s^2 }\] So the first thing is in units of meters that you're trying to add to m/s, so there's no way adding those two will give you units of acceleration because you can't add apples to oranges.

OpenStudy (kainui):

Perhaps try: \[a=\frac{ v_f-v_i }{ \Delta t }\] And remember, what is the final velocity? It's not 28 m/s. What's the speed when the goalie _stops_ it?

OpenStudy (anonymous):

I dont know

OpenStudy (kainui):

That's too bad.

OpenStudy (anonymous):

what grade level class is this? becuase technicaly speaking, the ball has an acceleration of 0 after it leaves the kicker and approaches the goalie because the velocity is constant between these two players. Constant velocity means acceleration is zero... so the 'average acceleration' between the two players, (for the duration of 1.6seconds) is zero. if the ball is traveling at a constant 28m/s

OpenStudy (anonymous):

It is grade level 8th

OpenStudy (anonymous):

Hello @DragonFlame, Well, if you copied the question without mistakes, then I would have to agree with @DemolisionWolf. Think about it, the ball traveled at constant velocity--constant velocity IS zero acceleration. Of course, there is always the possibility you mistyped the question, or that you didn't provide us with all the information.

OpenStudy (anonymous):

I think they just want the fact that the ball went from its initial velocity to zero velocity in the time of 1.6 s. This is an "average acceleration" in A CRUDE SENSE.